# SG-GS: Photo-realistic Animatable Human Avatars with Semantically-Guided Gaussian Splatting

### v1

@@ Introduction
where $\mathbf{V}\in\mathbb{R}^{M\times 3}$ and $M$ is the number of voxels. We then construct a 3D sparse U-Net by stacking a series of sparse convolutions with skip connections to aggregate local features. The sparse 3D U-Net takes $\mathbf{V}$ as input, denoted as $\mathbf{F}_{v}\in\mathbb{R}^{M\times C}$, where $C$ is the feature dimension.
To fuse the geometric and semantic information, we concatenate the voxel-based features $\mathbf{F}_{v}$ with the semantic point-based features $\mathcal{O}$, and then process them through an MLP layer:
$\Delta{(\mathcal{X}^{\prime},s^{\prime},r^{\prime})}=\text{MLP}(\text{Concat}(\mathbf{F}_{v},\mathcal{O})),$ (11)
where $\Delta{(\mathcal{X}^{\prime},s^{\prime},r^{\prime})}$ represents the final fused features. The deformed 3D Gaussian $\mathcal{G}_{o}$ is then deformed by $\Delta{(\mathcal{X}^{\prime},s^{\prime},r^{\prime})}$ following Eq. 4, 5, and 6.
Optimization
During densification, newly created 3D Gaussian points inherit semantic attributes from their parent nodes.
@@
$\mathcal{S}=\sum_{p\in\mathcal{N}}y_{p}\alpha_{p}\prod_{j=1}^{p-1}\left(1-\alpha_{j}\right),$ (12)
where $\mathcal{S}_{k}$ represents the 2D semantic labels of pixel $k$, derived from Gaussian point semantic attributes via $\alpha$-blending (Eq. 2). Here, $y_{p}$ denotes the semantic attribute of the 3D Gaussian point $p$, and $\alpha_{p}$ is the influence factor of this point in rendering pixels. Upon calculating these labels, we obtain the results $l^{s}_{i}$ and apply a BCE loss to regularize the rendered semantic label $l^{s}_{i}$ with semantic labels generated via SHA as follows:
@@
Semantic-guided density regularization. Fuzzy geometric shapes often appear in local structures on the human surface, particularly in high-frequency areas like clothing wrinkles and muscle textures (Wang et al. 2024). To improve the clarity and distribution of 3D Gaussians in these regions, we propose semantic-guided density regularization. We identify high-frequency nodes by assessing the average magnitude of structural differences between a selected node and all nodes within the same cluster. Nodes exhibiting the highest average magnitude of these differences are designated as high-frequency nodes.
@@
where $H_{m}$ is the high-frequency node in cluster $C_{m}$, $A_{i}$ is the basic attribute of 3D Gaussian points (color, opacity, etc.), $C_{m}$ represents the set of all points with semantic attribute $m$, and $d(\cdot,\cdot)$ is a dissimilarity measure between two points. To better capture and express these local structures of significant discrepancies, we perform densification operations on these 3D Gaussians, enhancing the local rendering granularity to focus on guiding the split and attribute optimization of Gaussian points in these areas.
@@
$\mathcal{L}_{neighborhood}=\frac{1}{|N|}\sum_{m\in N}\sum_{n\in N_{k}(m)}D_{\text{KL}}(p_{m}||p_{n}),$ (15)
where $N$ represents the total number of Gaussian points, $N_{k}(m)$ contains the $k$ nearest neighbors of 3D Gaussian point $m$ in 3D space, $p_{m}$ and $p_{n}$ represent the predicted probability distributions for point $m$ and its neighbor $n$, respectively, and $D_{\text{KL}}(p_{m}||p_{n})$ calculates the KL divergence between the predicted distributions of point $m$ and its neighbor $n$.
Loss function. Our full loss function consists of a RGB loss $\mathcal{L}_{rgb}$, a mask loss $\mathcal{L}_{mask}$, a skinning weight regularization loss $\mathcal{L}_{skin}$, the as-isometric-as-possible regularization loss $\mathcal{L}_{isopos}$ following (Qian et al. 2024), Semantic projection with 2D regularization $\mathcal{L}_{semantic}$, and Semantic-aware regularization with neighborhood consistency $\mathcal{L}_{neighborhood}$:
@@
Figure 2: Qualitative Comparison on ZJU-MoCap (Peng et al. 2021b). We show that our SG-GS can produce realistic details in both rendered images and geometry, while other approaches struggle to generate smooth details.
Figure 3: Qualitative Comparison on H36M (Ionescu et al. 2013). By utilizing semantic information within human body, our SG-GS preserves better anatomical structures of the human body, producing high-quality results
Figure 4: Ablation Study on Geometric and Semantic Feature Learning, which helps erase artifacts and learn fine details like cloth wrinkles and human face.
Figure 5: Ablation Study on semantic projection with 2D regularization, which enhances semantic accuracy.
Table 1: Quantitative Results on ZJU-MoCap (Peng et al. 2021b). SG-GS achieves state-of-the-art performance across every method. The best and the second best results are denoted by pink and yellow. Frames per second (FPS) is measured on an RTX 3090. LPIPS* = LPIPS $\times$ 1000.
@@
NeuralBody (Peng et al. 2021b) 29.07 0.962 52.29 1.5
Ani-NeRF (Peng et al. 2021a) 29.17 0.961 51.98 1.1
HumanNeRF (Weng et al. 2022) 30.24 0.968 31.73 0.3
MonoHuman (Yu et al. 2023) 29.38 0.964 37.51 0.1
InstantAvatar (Jiang et al. 2023) 29.73 0.938 64.41 4.2
3DGS-Avatar (Qian et al. 2024) 30.62 0.965 30.28 50
GauHuman (Hu et al. 2024b) 30.79 0.960 32.73 180
GoMAvatar (Wen et al. 2024) 30.37 0.969 32.53 43
@@
Table 2: Quantitative Results on H36M (Ionescu et al. 2013).
@@
NARF (Noguchi et al. 2021) 23.00 0.898 22.27 0.881
NeuralBody (Peng et al. 2021b) 22.89 0.896 23.09 0.891
Ani-NeRF (Peng et al. 2021a) 23.00 0.890 22.55 0.880
ARAH (Wang et al. 2022c) 24.79 0.918 23.42 0.896
3DGS-Avatar (Qian et al. 2024) 32.89 0.982 32.50 0.983
@@
Experiment
In this section, we first compare SG-GS with recent SOTA methods (Peng et al. 2021b, a; Weng et al. 2022; Wang et al. 2022c; Yu et al. 2023; Qian et al. 2024; Hu et al. 2024b, a), demonstrating that our SG-GS achieves superior rendering quality. We then systematically ablate each component of the proposed method, showing their effectiveness in better rendering performance. All models are trained on one single NVIDIA RTX 3090 GPU. For further details of implementation, please refer to the Supp.Mat.
Dataset
ZJU-MoCap (Peng et al. 2021b). It records multi-view videos with 21 cameras and collects human poses using the marker-less motion capture system. We select six sequences (377, 386, 387, 392, 393, 394) from this dataset to conduct experiments. We also follow the same training/test split as HumanNeRF (Weng et al. 2022) and 3DGS-Avatar (Qian et al. 2024), i.e., one camera is used for training, while the remaining cameras are used for evaluation.
H36M (Ionescu et al. 2013). It captures multi-view videos using four cameras and collects human poses with a marker-based motion capture system. It includes multiple subjects performing complex actions. We select representative actions, split the videos into training and test frames, following ARAH (Wang et al. 2022c), and perform experiments on sequences (S1, S5, S6, S7, S8, S9, S11). Three cameras are used for training and the remaining is selected for test.
Comparison with State-of-the-art Methods
We conduct comparative experiments against various state-of-the-art (SOTA) methods for human avatars, including NeRF-based methods such as NeuralBody (Peng et al. 2021b), Ani-NeRF (Peng et al. 2021a), HumanNeRF (Weng et al. 2022), and MonoHuman (Yu et al. 2023), as well as 3DGS-based methods such as 3DGS-Avatar (Qian et al. 2024), GauHuman (Hu et al. 2024b), and GoMAvatar (Wen et al. 2024), under a monocular setup on ZJU-MoCap (Peng et al. 2021b). In Tab.1, we evaluate the reconstruction quality using three different metrics: PSNR, SSIM, and LPIPS. Thanks to the LBS weight field and deformation field learned in HumanNeRF(Weng et al. 2022), 3DGS-Avatar (Qian et al. 2024), and GauHuman (Hu et al. 2024b), these methods achieve comparable visualization results. In comparison, our proposed SG-GS achieves good performance in terms of PSNR and SSIM while significantly outperforming existing methods on LPIPS. Existing researches (Park et al. 2021; Yang et al. 2024a) have reached a consensus that LPIPS provides more meaningful insights compared to the other metrics, given the challenges of reproducing exact ground-truth appearances for novel views.
As shown in Fig. 2, our SG-GS method preserves sharper details compared to other methods. Notably, our approach excels at capturing fine details in challenging areas such as clothing, where reconstruction is typically more difficult due to intricate textures. By preserving these finer details, our method provides a more realistic and detailed reconstruction of clothing and other complex surfaces, significantly improving the overall quality and fidelity of the 3D human avatars.
In addition, we also evaluate our SG-GS using the H36M (Ionescu et al. 2013) dataset. We report the quantitative results against NeRF-based methods such as NARF (Noguchi et al. 2021), NeuralBody (Peng et al. 2021b), Ani-NeRF (Peng et al. 2021a), and ARAH (Wang et al. 2022c), as well as 3DGS-based methods such as 3DGS-Avatar (Qian et al. 2024) in Tab.2. Our model outperforms both established NeRF-based methods and 3DGS-based methods.
As shown in Fig.3, due to the use of semantic information within human body, our SG-GS achieves better reconstruction of edge areas and preserves anatomical structures of the human body.
Table 3: Ablation Study on ZJU-MoCap (Peng et al. 2021b). The proposed model achieves the lowest LPIPS, demonstrating the effectiveness of all components.
@@
Ablation Study
@@
In this section, we evaluate the effectiveness of our proposed modules through ablation experiments on the ZJU-MoCap dataset. These modules include Geometric and Semantic Feature Learning (geo), semantic projection with 2D regularization ($\mathcal{L}_{semantic}$), semantic-guided density regularization (sgd), and semantic-aware regularization with Neighborhood Consistency ($\mathcal{L}_{neighborhood}$). The average metrics over 6 sequences are shown in Table 3.
@@
Geometric and Semantic Feature Learning. This module is designed to learn the associations between body geometry and semantics for 3D deformation. As shown in Table 3, it significantly enhances rendering performance. Though it slightly increase inference time, the notable performance improvement justifies this additional cost. A qualitative comparison in Fig. 4 further proves that Geometric and Semantic Feature Learning reduces artifacts and captures fine details like cloth wrinkles and facial features.
Semantic Projection with 2D Regularization. This part utilizes semantic labels generated by SHA to supervise the semantic attributes of 3D Gaussians. As shown in Fig. 5, semantic projection with 2D regularization substantially improves the semantic accuracy of 3D Gaussians. The results in Table 3 illustrate the critical role of semantic information in creating photo-realistic animatable human avatars.
@@
Semantic-Guided Density Regularization and Semantic-Aware Regularization with Neighborhood Consistency. Semantic-guided density regularization enhances rendering quality by optimizing Gaussian density in areas with high discrepancy, while semantic-aware regularization with neighborhood consistency ensures that nearby Gaussians exhibit coherent semantic attributes, thus improving 3D semantic consistency. The improvements in rendering quality are validated by the results in Table 3.
Conclusion
@@
In this paper, we propose SG-GS, which uses semantics-embedded 3D Gaussians to reconstruct photo-realistic animatable human avatars. SG-GS first integrates a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to deform photo-realistic animatable human avatars. SG-GS then leverages SMPL’s human body semantic priors to acquire human body semantic labels, which are used to guide optimization of Gaussian’s semantic attribute. To utilize local geometric and semantic information effectively, we propose a 3D geometry- and semantic-aware network to learn body geometric and semantic associations and integrate them into the 3D deformation. We further implement three key strategies to enhance semantic accuracy and render quality: semantic projection with 2D regularization, semantic-guided density regularization, and semantic-aware regularization with neighborhood consistency. Extensive experiments demonstrate that SG-GS outperforms SOTA methods in creating photo-realistic animatable avatars, further validating our hypothesis that integrating semantic priors enhances fine-detail reconstruction. We hope that our method will foster further research in high-quality clothed human avatar synthesis from monocular views.
@@
Limitations. 1). SG-GS lacks the capability to extract 3D meshes. Developing a method to extract meshes from 3D Gaussians is an important direction for future research. 2). Geometric and Semantic Feature Learning employs a 3D sparse U-Net, which is computationally intensive and may increase training and inference time to some extent.

### v2

@@ 4.3 Topological and Geometric Feature Learning
where $\mathbf{V}\in\mathbb{R}^{M\times 3}$ and $M$ is the number of voxels. We then construct a 3D sparse U-Net by stacking a series of sparse convolutions with skip connections to aggregate local features. The sparse 3D U-Net $f_{\theta_{unet}}$ takes $\mathbf{V}$ and the semantic point-based features $\mathcal{O}$ as input, and outputs topological and geometric features $\mathbf{F}_{v}$:
$\mathbf{F}_{v}=f_{\theta_{unet}}(\mathbf{V};\mathcal{O}).$ (10)
We process the feature $\mathbf{F}_{v}$, the position $\mathcal{X}_{d}$ of the deformed Gaussians $\mathcal{G}_{o}$, and pose latent code $\mathcal{Z}_{p}$ in Eq. 2 through an fusion network $f_{\theta_{sr}}$:
$\Delta{(\mathcal{X}^{\prime},s^{\prime},r^{\prime})}=f_{\theta_{sr}}(\mathbf{F}_{v};\mathcal{X}_{d};\mathcal{Z}_{p}),$ (11)
where $\Delta{(\mathcal{X}^{\prime},s^{\prime},r^{\prime})}$ represents the final fused features. The deformed 3D Gaussian $\mathcal{G}_{o}$ is then deformed by $\Delta{(\mathcal{X}^{\prime},s^{\prime},r^{\prime})}$ following Eq. 3, 4, and 5.
4.4 Optimization
Unlike random initialization or Structure-from-Motion (SfM) initialization for Gaussian point clouds, we directly sample 6,890 points from the SMPL model [22] as our initial point cloud. Each Gaussian is then assigned semantic attributes based on the SMPL model’s predefined semantic labels.During densification, newly created 3D Gaussian points inherit semantic attributes from their parent nodes.
@@ 4.4 Optimization
$\mathcal{S}=\sum_{g\in\mathcal{N}}\mathcal{O}_{g}\alpha_{g}\prod_{j=1}^{g-1}\left(1-\alpha_{j}\right),$ (12)
where $\mathcal{S}_{k}$ represents the 2D semantic labels of pixel $k$, derived from Gaussian point semantic attributes via $\alpha$-blending (Eq. 1). Here, $\mathcal{O}_{g}$ denotes the semantic attribute of the 3D Gaussian point $g$, and $\alpha_{g}$ is the influence factor of this point in rendering pixels. Upon calculating these labels, we obtain the results $l^{s}_{i}$ and apply a BCE loss to regularize the rendered semantic label $l^{s}_{i}$ with semantic labels generated via SHA as follows:
@@
Semantic-guided density regularization. Fuzzy geometric shapes often appear in local structures on the human surface, particularly in high-frequency areas like clothing wrinkles and muscle textures [37]. To improve the clarity and distribution of 3D Gaussians in these regions, we propose semantic-guided density regularization. We identify high-frequency nodes by assessing the average magnitude of structural differences between a selected node and all nodes within the same cluster. Nodes exhibiting the highest average magnitude of these differences are designated as high-frequency nodes.
@@
where $H_{m}$ is the high-frequency node in cluster $C_{m}$, $A_{i}$ is the basic attribute of 3D Gaussian points (color, opacity, etc.), $C_{m}$ represents the set of all points with semantic attribute $m$, $C_{m}\setminus\{i\}$ denotes the set of elements in $C_{m}$ excluding the element $i$, and $d(\cdot,\cdot)$ is a dissimilarity measure between two points. To better capture and express these local structures of significant discrepancies, we perform densification operations on these 3D Gaussians, enhancing the local rendering granularity to focus on guiding the split and attribute optimization of Gaussian points in these areas.
@@
$\mathcal{L}_{neighborhood}=\frac{1}{|N|}\sum_{m\in N}\sum_{n\in N_{k}(m)}D_{\text{KL}}(\mathcal{O}_{m}||\mathcal{O}_{n}),$ (15)
where $N$ represents the total number of Gaussian points, $N_{k}(m)$ contains the $k$ nearest neighbors of 3D Gaussian point $m$ in 3D space, $\mathcal{O}_{m}$ and $\mathcal{O}_{n}$ represent the predicted semantic attribute for point $m$ and its neighbor $n$, respectively, and $D_{\text{KL}}(q_{m}||q_{n})$ calculates the KL divergence between the predicted distributions of point $m$ and its neighbor $n$.
Figure 3: Qualitative Comparison on ZJU-MoCap [31]. We show that our SG-GS can produce realistic details in both rendered images and geometry, while other approaches struggle to generate smooth details.
Loss function. Our full loss function consists of a RGB loss $\mathcal{L}_{rgb}$, a mask loss $\mathcal{L}_{mask}$, a skinning weight regularization loss $\mathcal{L}_{skin}$, the as-isometric-as-possible regularization loss $\mathcal{L}_{isopos}$ following [32], Semantic projection with 2D regularization $\mathcal{L}_{semantic}$, and Semantic-aware regularization with neighborhood consistency $\mathcal{L}_{neighborhood}$:
@@
Table 1: Quantitative Results on ZJU-MoCap [31]. SG-GS achieves state-of-the-art performance across every method. The best and the second best results are denoted by pink and yellow. Frames per second (FPS) is measured on an RTX 3090. LPIPS* = LPIPS $\times$ 1000.
@@
NeuralBody [31] 29.07 0.962 52.29 1.5
Ani-NeRF [30] 29.17 0.961 51.98 1.1
HumanNeRF [40] 30.24 0.968 31.73 0.3
MonoHuman [46] 29.38 0.964 37.51 0.1
DVA [33] 29.45 0.956 37.74 17
InstantAvatar [12] 29.73 0.938 64.41 4.2
3DGS-Avatar [32] 30.62 0.965 30.28 50
GauHuman [10] 30.79 0.960 32.73 180
GoMAvatar [39] 30.37 0.969 32.53 43
@@
Table 2: Quantitative Results on H36M [11]. Our SG-GS still achieves superior performance compared to state-of-the-art methods on both training poses and novel poses.
@@
NARF [29] 23.00 0.898 22.27 0.881
NeuralBody [31] 22.89 0.896 23.09 0.891
Ani-NeRF [30] 23.00 0.890 22.55 0.880
ARAH [38] 24.79 0.918 23.42 0.896
3DGS-Avatar [32] 32.89 0.982 32.50 0.983
@@
Figure 4: Qualitative Comparison on H36M [11]. By utilizing semantic information within human body, our SG-GS preserves better anatomical structures of the human body, producing high-quality results
Table 3: Ablation Study on ZJU-MoCap [31]. The proposed model achieves the lowest LPIPS, demonstrating the effectiveness of all components.
@@
5 Experiment
In this section, we first compare SG-GS with recent SOTA methods [31, 30, 40, 38, 46, 32, 10, 9], demonstrating that our SG-GS achieves superior rendering quality. We then systematically ablate each component of the proposed method, showing their effectiveness in better rendering performance. All models are trained on one single NVIDIA RTX 3090 GPU. For further details of implementation, please refer to the Supp.Mat.
5.1 Dataset
ZJU-MoCap [31]. It records multi-view videos with 21 cameras and collects human poses using the marker-less motion capture system. We select six sequences (377, 386, 387, 392, 393, 394) from this dataset to conduct experiments. We also follow the same training/test split following [40, 32], i.e., one camera is used for training, while the remaining cameras are used for evaluation.
H36M [11]. It captures multi-view videos using four cameras and collects human poses with a marker-based motion capture system. It includes multiple subjects performing complex actions. We select representative actions, split the videos into training and test frames, following ARAH [38], and perform experiments on sequences (S1, S5, S6, S7, S8, S9, S11). Three cameras are used for training and the remaining is selected for test.
5.2 Comparison with State-of-the-art Methods
We conduct comparative experiments against various state-of-the-art (SOTA) methods for human avatars, including NeRF-based methods such as NeuralBody [31], Ani-NeRF [30], HumanNeRF [40], and MonoHuman [46], as well as 3DGS-based methods such as 3DGS-Avatar [32], GauHuman [10], and GoMAvatar [39], under a monocular setup on ZJU-MoCap [31]. In Table 1, we evaluate the reconstruction quality using three different metrics: PSNR, SSIM, and LPIPS. Thanks to the LBS weight field and deformation field learned in HumanNeRF[40], 3DGS-Avatar [32], and GauHuman [10], these methods achieve comparable visualization results. In comparison, our proposed SG-GS achieves good performance in terms of PSNR and SSIM while significantly outperforming existing methods on LPIPS. Existing researches [43, 32] reach a consensus that LPIPS provides more meaningful insights compared to the other metrics, given the challenges of reproducing exact ground-truth appearances for novel views.
As shown in Fig. 3, our SG-GS method preserves sharper details compared to other methods. Notably, our approach excels at capturing fine details in challenging areas such as clothing, where reconstruction is typically more difficult due to intricate textures. By preserving these finer details, our method provides a more realistic and detailed reconstruction of clothing and other complex surfaces, significantly improving the overall quality and fidelity of the 3D human avatars. Please see our project website videos and supplementary material for more video visualization.
In addition, we also evaluate our SG-GS using the H36M [11] dataset. We report the quantitative results against NeRF-based methods such as NARF [29], NeuralBody [31], Ani-NeRF [30], and ARAH [38], as well as 3DGS-based methods such as 3DGS-Avatar [32] in Table 2. Our model outperforms both established NeRF-based methods and 3DGS-based methods. As shown in Fig. 4, due to the use of semantic information within human body, our SG-GS achieves better reconstruction of edge areas and preserves anatomical structures of the human body.
Figure 5: Ablation Study on Geometric and Semantic Feature Learning, which helps erase artifacts and learn fine details like cloth wrinkles and human face under novel views.
Figure 6: Ablation Study on semantic projection with 2D regularization, which enhances semantic accuracy. During pruning, most Gaussians are removed, leaving the remaining ones to default to torso semantics without our semantic supervision.
Figure 7: Ablation Study on semantic projection with 2D regularization, which keeps the topological consistency of the human body under novel poses.
5.3 Ablation Study
@@
In this section, we evaluate the effectiveness of our proposed modules through ablation experiments on the ZJU-MoCap [31] dataset. The average metrics over 6 sequences are shown in Table 3.
@@
Topological and Geometric Feature Learning. As shown in Table 3, the proposed module significantly (topo-geo) enhances rendering performance. Though it slightly increase inference time, the notable performance improvement justifies this additional cost. A qualitative comparison in Fig. 5 further proves that Topological and Geometric Feature Learning maintains anatomical coherence during motion and preserving fine details. We also conduct an experiment replacing the sparse 3D U-Net with an MLP (mlp in Table 3), which demonstrates point-level MLP is limited by a small receptive field, restricting capability to capture the local geometric and topological features.
Semantic Projection with 2D Regularization. This part utilizes semantic labels generated by SHA to supervise the semantic attributes of 3D Gaussians ($\mathcal{L}_{semantic}$). As shown in Fig. 6, semantic projection with 2D regularization substantially improves the semantic accuracy of 3D Gaussians. At the start of training, Gaussians are neither densified nor pruned [32], allowing their scale to grow. During pruning phase, most Gaussians are removed. As a result, most remaining Gaussians default to torso semantics without supervision. The results ($\mathcal{L}_{semantic}$ in Table 3) highlight the critical role of semantic information. This demonstrates that while the sparse 3D U-Net (introduced in Section 4.3) can capture geometric features with noisy semantic information and improve rendering quality, it still requires accurate semantic data to learn the topology to keep anatomical coherence of the human body, as shown in Fig. 7.
@@
Semantic-Guided Density Regularization and Semantic-Aware Regularization with Neighborhood Consistency. Semantic-guided density regularization (sgd) enhances rendering quality by optimizing Gaussian density in areas with high discrepancy, while semantic-aware regularization with neighborhood consistency ($\mathcal{L}_{neighborhood}$) ensures that nearby Gaussians exhibit coherent semantic attributes, thus improving 3D semantic consistency. The improvements in rendering quality are validated by the results in Table 3.
6 Conclusion
@@
In this paper, we propose SG-GS, which uses semantics-embedded 3D Gaussians to reconstruct photo-realistic human avatars. SG-GS first integrates a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to deform human avatars. SG-GS then leverages SMPL’s human body semantic priors to acquire human body semantic labels, which are used to guide optimization of Gaussian’s semantic attribute. We also propose a 3D topology- and geometry-aware network to learn body geometric and topological associations and integrate them into the 3D deformation. We further implement three key strategies to enhance semantic accuracy and render quality: semantic projection with 2D regularization, semantic-guided density regularization, and semantic-aware regularization with neighborhood consistency. Extensive experiments demonstrate that SG-GS outperforms SOTA methods in creating photo-realistic avatars, further validating our hypothesis that integrating semantic priors enhances fine-detail reconstruction. We hope that our method will foster further research in high-quality clothed human avatar synthesis from monocular views.
@@
Limitations. 1). SG-GS lacks the capability to extract 3D meshes. Developing a method to extract meshes from 3D Gaussians is an important direction for future research. 2). Topological and Geometric Feature Learning employs a sparse 3D U-Net, which is computationally intensive and may increase training and inference time to some extent.